OpenStudy (anonymous):

Use the grouping method to factor the polynomial below completely. 5x5 + 30x3 + 25x

OpenStudy (hoblos):

5x5 + 30x3 + 25x = 5x(x4+6x2+5) = 5x(x2+5)(x2+1)

OpenStudy (anonymous):

5x^5 + 30x^3 + 25x =5x^5 + 5x^3 +25x^3 +25x =5x^3 (x^2 +1) + 25x(x^2+1) =(x^2+1)(5x^3 +25x) =(x^2+1)(5x)(x^2+5) =5x(x^2+1)(x^2+5)